1. A semicircular window of radius 15 inches is to be laminated with a sunblock coating that costs $0.85 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use p = 3.14. Hint: Area of Circle: A = prē)
Thank you for your help. If you could lay it our for me, I'd appreciate it.
AD

Answer by checkley77(12569)

area=pi*r^2
are=3.14*15^2/2=353.25 in^2.
352.25*.85=$300.26 cost to coat.